OpenStudy (gabbyalicorn):

A bag contains 8 black marbles, 16 white marbles, 14 red marbles, and 10 green marbles. A marble is drawn at random from the bag. What is the probability that the marble drawn will not be white?

OpenStudy (hockeychick23):

8+16+14+10= ?

OpenStudy (gabbyalicorn):

A 1/3 B 2/3 C 3/4 D 15/16

OpenStudy (gabbyalicorn):

yes i was going to do that @hockeychick23

OpenStudy (gabbyalicorn):

16/48? but then simplyfy

OpenStudy (anonymous):

divide by 4

OpenStudy (anonymous):

4/12

OpenStudy (hockeychick23):

not quite, you want to know the probability of drawing one that isn't white

OpenStudy (hockeychick23):

so 48-16= ?

OpenStudy (gabbyalicorn):

8/24 = 4/12 = 2/6= 1/2 ?

OpenStudy (hockeychick23):

32/48= 4/6= 2/3

OpenStudy (gabbyalicorn):

oh :/ darn

OpenStudy (gabbyalicorn):

Thank you @hockeychick23 for your help!

OpenStudy (hockeychick23):

sure!(: